OpenStudy (aravindg):

find limit x tends to 4 (x-4)/(x^2-x-12)

OpenStudy (aravindg):

can u giv steps???

OpenStudy (mimi_x3):

(x-4)/(x-4)(x+3) = 1/(x+3) x -> 4 = 1/ (4+3) = 1/7

OpenStudy (aravindg):

thx a lot 4 more plz wait

OpenStudy (anonymous):

(x-4)/(x^2-x-12) simplifies to (x-4)/(x+3)(x-4) = 1 / (x+3) as x approaches 4 it tends to 1/ (4+3) = 1/7

OpenStudy (aravindg):

lim x tend to 3 x^3 -27/x^2-9

OpenStudy (aravindg):

.........fast plzz

OpenStudy (mimi_x3):

relax mans it takes time , try doing it yurself :P

OpenStudy (anonymous):

yeah aravind what's wrong with you it takes time

OpenStudy (mimi_x3):

= 3(x^3 -9) / (x^2-9) = 3(x-3)(x^2+3x+9)/(x-3)(x+3) = 3(x^2 + 3x +9)/(x+3) = 21/2 ( i have a feeling that its wrong though xD

OpenStudy (anonymous):

no it is correct mimi just remove the 3 it must be some confusion otherwise you solution is perfect 9/2 is the answer

OpenStudy (mimi_x3):

remove the 3 ?

OpenStudy (anonymous):

\[\frac{x^3-3^3}{x^2-9}\] where x tend to 3

OpenStudy (mimi_x3):

yu jst eliminate the 3 out of no where ? lols , im confused xD

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3}\frac{x^3-3^3}{x^2-3^2}\]This is the question/problem \[\lim_{x \rightarrow 3} \frac{(x-3)(x^2+3x+9)}{(x-3)(x+3)}\] \[when,\:\:\:x \rightarrow 3\] \[\frac{9}{2}\]

OpenStudy (mimi_x3):

ohh yeah , my bad i read the question wrongly i thought it was 3x^3 -27/x^2-9 lols , wonder why i got confused lmao